OpenStudy (anonymous):

\[at what piont is the function f \left( x \right)=[x -3]+\left[ x -3 \right],-4\le x \le4\]

OpenStudy (anonymous):

\[at what piont is the function f \left( x \right)=[x -3]+\left[ x -3 \right],-4\le x \le4\]